# Liu He1, Yizhi Song1, Hejun Huang2, Daniel Aliaga1, Xin Zhou2

### v1

@@ Introduction
Fig. 4A to D show visual comparisons on four text descriptions with increasing complexity. Although Pika and Gen-2 (Runaway 2024) show natural and realistic lighting, they fail to synthesize highly dynamic motions and only show minor motion flows between key frames. Given complex descriptions, Gen-2 misunderstands camera control prompts as the main character (Fig. 4C and D), and Luma generates a meaningless character in Fig. 4D. All baselines are prone to visual artifacts (especially in arms and garments) and motion inconsistency. In contrast, our videos have higher motion magnitude, and better camera control and temporal consistency.
@@
In addition, we conduct a user study (105 users, ranging in age from 21 to 40) comparing our results to the three baseline models. We randomly sample 10 videos from our generated videos and those by competitors using the same video description input. For each question, the user will read a video description, then watch four videos (one video from each model) in random order. The user is asked to select one best choice for each of the four criteria: prompt following, character motion, camera motion, and overall quality. We report the win rate from each model and for each criteria in Tab. 2. The table shows that our method significantly outperforms all baseline models. It also highlights the superiority of our model to follow the character motion and camera control instructions embedded in the video text description. Additional details are provided in the Supplemental.
Method Prompt Following $\uparrow$ Character Motion $\uparrow$ Camera Motion $\uparrow$ Overall Quality $\uparrow$ Gen-2 13.77 16.69 10.23 17.49 Pika 4.72 9.03 10.64 7.23 Luma 27.13 26.93 24.90 31.58 Ours 54.37 47.35 54.23 43.70
Table 2: User Study. Based on the side-by-side comparisons of our generated video to other three videos generated by the baseline models, our model outperforms other models by much higher win rate (%) in all four criteria.
Ablation Studies
RAG Contributions.
@@
LLM/VLM Selection.
We alternatively tested Claude-3-Sonnet (Claude 2024) as the backbone for all agents and also alternatively used LLaMa-3-70B-Instruct (Dubey et al. 2024) as the backbone of LLM-Director/Programmer. However, in all cases, using GPT-4V excels in coding and visual understanding. For example, using GPT-4V reduces in average 34.79% reviewing iterations compared to using other LLM/VLMs.
Limitations
@@
Conclusion
End-to-end video generation has been struggling with physical correctness, camera control, and temporal consistency. While procedural approaches require tedious manual crafting, our method successfully leverages LLM/VLM agents and 3D engines (e.g. Blender) for automatic synthetic video generation. We design collaborative generation, visual reviewing, and library updating phases as part of a robust and efficient pipeline, and use RAG with video making and 3D engine knowledge for further improvement. Our approach outperforms 3 commercial video generation models in multiple metrics and in our user study regarding prompt following, character motion, camera motion, and visual fidelity.
Moving forward, our method may provide a synthetic video prior for foundation model training, video-to-video generation, and virtual storyboarding.
Supplemental Materials
Below is a summary of the contents in each section of this supplemental material:
  • Showcase of more generated videos by our framework.
  • Video-to-video style transferring by implementing diffusion model per key frame.
  • Example text/multimodal prompts given as input to our three agents.
  • Examples of library composing and updating, and composing scripts calling library by LLM-Programmer.
  • More details about the user study.
Generated Videos
In Supplemental Fig 1, we provide seven additional video clips as in Fig.1 (main paper). The generated videos show a good ability of instruction-following, consistent character motion, and smooth camera controlling.
@@ Conclusion
Video-to-Video Style Transfer
In Supplemental Fig. 2, we show video-to-video style transfer by applying the diffusion-based model AnimateDiff (Guo et al. 2023) and ControlNet (Zhang, Rao, and Agrawala 2023) (given both canny edge and depth) to each key frame output of our method. It transfers our input images into the altered animation style. The converted video inherits consistent character motion and smooth camera motion from the video generated by our method. It shows the potential of using our generated videos as prior for style-varying video generation.
Figure 2: Video-to-Video Style Transfer. We use AnimateDiff (Guo et al. 2023) and ControlNet (Zhang, Rao, and Agrawala 2023) (given both canny edge and depth) for each key frame output by our method. The converted video (lower row) inherits consistent character motion and smooth camera motion from the video generated by our method (upper row). It shows the potential of using our generated videos as prior for diversely-styled video generation. Note that the color inconsistency across key frames results from the randomness in the diffusion model.
Prompts for Agents
@@
Library Composing and Updating
@@
User Study
In addition to the information about the user study provided in the main paper (Qualitative Comparisons Section), in Fig. 10 we show an example user study page containing videos generated by our method and three commercial baselines (placed in random order). The user will see all videos and corresponding video descriptions. Then, the user will select a best choice for each of the four criteria: prompt following, character motion, camera motion, and overall quality. Each user will see a total of 10 pages as Fig. 10. Note that we cover the upper right and lower right corner of all videos using same-sized black boxes in order to prevent the watermarks of the baseline methods from being visible.
Figure 10: User Study Page. An example user study page containing videos generated by our method and three commercial baselines (placed in random order). Upper right and lower right corner of all videos are blocked by the same-sized black boxes to hide logos.

### v2

@@ 4.4 Qualitative Comparisons
Fig. 5 shows visual comparisons on text descriptions. We provide additional visual results in the appendix Fig. 2 as C and D. Although Pika and Gen-2 [46] show natural and realistic lighting, they fail to synthesize highly dynamic motions and only show minor motion flows between key frames. Given complex descriptions, Gen-2 misunderstands camera control prompts as the main character (in C), and the results generated by Luma fail to follow the character motion prompts (in A and B). Although CogVideo-X and WAN-2.1 produce high-resolution and more realistic videos, they still struggle with large motion patterns, leading to significant blurriness. All baseline models are prone to visual artifacts, particularly in arms and garments, and suffer from motion inconsistency. In contrast, our approach achieves greater motion magnitude, precise camera control, and superior temporal consistency.
User Study
@@
In addition, we conducted a user study to compare the seven baseline video generation methods with ours. 30 users participated in our study. For each video description, we randomly sample 2 videos from our generated videos and from each of the baseline methods (a total of 16 videos). The user is first presented with a video description and then watches the corresponding video. The order of the videos is random and the method used is hidden from the user. The user is asked to rate the video on a scale of 1 to 5 for each of four criteria: (1) prompt following, (2) character motion stability, (3) camera motion quality, and (4) overall quality. We report the average score for each model and for each criteria in Tab. 2. The table shows that our method significantly outperforms all baseline models. It also highlights the superiority of our model in following the character motion and camera control instructions embedded in the video text description. More details are provided in the Supplementary Fig. 11.
Figure 6: Motion and Camera Analysis. We prompt the system with “A girl waving her hands with camera tilt up.” Compared to existing generation frameworks, our approach produces more diverse character motions and dynamic camera movements. Additionally, current methods often struggle with large and rapid motion patterns, leading to noticeable blurring artifacts.
Analysis
Compared to existing end-to-end text-to-video frameworks, our LLM-based approach demonstrates significantly improved character motion quality and more dynamic camera control. As shown in Fig. 6, our method does present motion blur, which is a common issue in traditional models when handling large or fast movements. By leveraging the reasoning capabilities of LLMs and agent-based framwork, our system generates a wider variety of motion patterns, leading to more natural and expressive character animation. Additionally, our approach enhances camera motion control, enabling smoother and more intentional movements. These improvements collectively contribute to a more visually coherent and aesthetically pleasing motion synthesis, making our method a strong alternative to conventional text-to-video generation techniques.
4.5 Ablation Studies
RAG Contributions
@@ 4.5 Ablation Studies
LLM/VLM Selection
We tested Claude-3-Sonnet [11] as the alternative backbone for all agents and also used LLaMa-3-70B-Instruct [13] as the alternative backbone of LLM-Director/Programmer. However, in all cases, using GPT-4V excels in coding and visual understanding. For example, using GPT-4V reduces in average 34.79% reviewing iterations compared to using other LLM/VLMs.
4.6 Limitations
@@
5 Conclusion
End-to-end video generation has been struggling with physical correctness, camera control, and temporal consistency. While procedural approaches require tedious manual crafting, our method successfully leverages LLM/VLM agents and 3D engines (e.g. Blender) for automatic synthetic video generation. We design collaborative generation, visual reviewing, and library updating phases as part of a robust and efficient pipeline, and use RAG with video making and 3D engine knowledge for further improvement. Our approach outperforms seven commercial or open-sourced video generation models in multiple metrics, and for the ability of prompt following, character motion, camera motion, and visual fidelity as per our user study.
Supplementary Material
1 Generated Videos
In Supplemental Fig. 1, we provide seven additional video clips as in Fig.1 (main paper). The generated videos show a good ability of instruction-following, consistent character motion, and smooth camera controlling.
@@ 1 Generated Videos
2 Additional Qualitative Comparisons
This part is a further extension of the qualitative comparison in the main paper. The visual comparisons are shown in Fig. 2. The results of the baseline models in C suggest that prompts of a fixed camera always lead to more static or less dynamic motions of the objects. Our agent-based framework does not present this issue. For the example in D, our method presents higher motion dynamics over existing frameworks.
Figure 2: More Qualitative Comparisons. Our method excels in handling character/object motions (highlighted in blue) and complex camera moves (highlighted in orange).
3 Video-to-Video Style Transfer
In Fig. 3, we show video-to-video style transfer by applying the diffusion-based model AnimateDiff [17] and ControlNet [72] (given both canny edge and depth) to each key frame output of our method. It transfers our input images into the altered animation style. The converted video inherits consistent character motion and smooth camera motion from the video generated by our method. It shows the potential of using our generated videos as prior for style-varying video generation.
Figure 3: Video-to-Video Style Transfer. We use AnimateDiff [17] and ControlNet [72] (given both canny edge and depth) for each key frame output by our method. The converted video (lower row) inherits consistent character motion and smooth camera motion from the video generated by our method (upper row). It shows the potential of using our generated videos as prior for diversely-styled video generation. Note that the color inconsistency across key frames results from the randomness in the diffusion model.
4 Prompts for Agents
@@
5 Library Composing and Updating
@@
6 User Study
In addition to the information about the user study provided in the main paper (Qualitative Comparisons Section), in Fig. 11 we show an example user study page containing videos generated by our method and seven baselines. The user will see all videos along with their descriptions and rate each video based on four criteria: prompt following, motion quality, camera move, and overall quality, on a scale of 1 to 5. Each user will evaluate a total of 10 pages similar to Fig. 11. Note that we cover the upper right and lower right corners of all videos with same-sized black boxes to obscure baseline method watermarks.
Figure 11: User Study Page. An example user study page for evaluating video generation quality. Participants rate videos based on four metrics: Prompt Following, Motion Quality, Camera Move, and Overall Quality, on a scale of 1 to 5. Videos from our method and 7 other baselines are randomly placed. Black boxes hide logos in the upper right and lower right corners.